OpenStudy (anonymous):

Determine whether the series 2/((n^2)-1) is convergent or divergent.

OpenStudy (tkhunny):

Telescoping? \[\frac{2}{n^{2}-1} = \frac{1}{n-1}-\frac{1}{n+1}\] for n>1